OpenStudy (anonymous):

2. As a new element is heated, it’s temperature can be modeled with the function T(x)= √100𝑥2−80𝑥+3500 where x represents the times, in minutes, from when the element is placed on top of a Bunsen burner, and T(x) is the temperature in degrees Celsius. When will the temperature be 100 degrees C.

OpenStudy (anonymous):

@Michele_Laino please help

OpenStudy (michele_laino):

we have to find the time \(x_0\) such that the subsequent condition holds: \[\huge T\left( {{x_0}} \right) = 100\]

OpenStudy (michele_laino):

so, we have to solve this equation: \[\huge \sqrt {100} {x^2} - 80x + 3500 = 100\]

OpenStudy (michele_laino):

oops... \[\huge \sqrt {100} x_0^2 - 80{x_0} + 3500 = 100\] being: \[\huge T\left( {{x_0}} \right) = \sqrt {100} x_0^2 - 80{x_0} + 3500\]

OpenStudy (michele_laino):

please solve such quadratic equation